### v1

# Mixture of Experts Using Tensor Products

Abstract
In multi-task learning, the conventional approach involves training a model on multiple tasks simultaneously. However, the training signals from different tasks can interfere with one another, potentially leading to negative transfer. To mitigate this, we propose a novel modular language model (TensorPoly), equipped with an inventory of latent modules and featuring a dynamic routing function that allocates modules to each task. Specifically, we reparameterize Low-Rank Adaptation (LoRA) by employing an entangled tensor through the use of tensor product operations, and call the resulting approach TLoRA. We tailor two innovative routing functions according to the granularity of the routing function: TensorPoly-I which directs to each rank within the entangled tensor while TensorPoly-II offers a finer-grained routing approach targeting each order of the entangled tensor. The experimental results from the multi-task T0-benchmark demonstrate that our modular LMs TensorPoly surpass the strong dense approaches, highlighting the potential of modular language models to mitigate negative inference in multi-task learning and deliver superior outcomes.
1 Introduction
Recently, the de facto paradigm for natural language understanding (NLU) tasks has centered on leveraging large language models (he2021towards) are pre-trained on a vast corpus of unlabelled data and subsequently fine-tuned for specific tasks (qiu2020pre; ye2021crossfit). While this approach has significantly advanced the field, it often requires substantial computational resources and may not efficiently transfer knowledge across diverse tasks. In addition, fine-tuning tasks independently can lead to negative transfer, where the lack of shared information across tasks, fails to achieve compositional generalization (poly).
To address the aforementioned issues, there have, roughly, been two general lines of research. The first line is to mitigate the computation and memory issue, several lightweight alternatives known as parameter-efficient finetuning (PEFT) have been proposed to update only a small number of extra parameters while keeping most pre-trained parameters frozen (houlsby2019parameter; li2021prefix; lora). However, these solutions need to train each adapter for each task, which does not take into account that test tasks may require solving different combinations of sub-problems compared to training tasks (vu2020exploring), thus failing to achieve compositional generalization (rosenbaum2019routing; ponti2021inductive). Moreover, specializing the model toward different tasks during training can yield gradients that are poorly aligned, resulting in negative transfer (wang2020gradient; lialin2023scaling).
The second line is to facilitate information sharing across multiple tasks using multi-task learning (MTL) approaches (caruana1997multitask; zhang2021survey; liu2019multi), which simultaneously trains the model on several tasks, allowing it to learn shared representations that benefit all tasks involved. However, MTL necessitates access to all training tasks during the training phase, meaning that incorporating new tasks requires retraining the model from scratch. This requirement significantly increases the computational burden and limits the flexibility of the model to adapt to new tasks efficiently.
A promising approach to address the above issues is the adoption of modular deep learning (pfeiffer2023modular), where modules are typically implemented as parameter-efficient approaches (lora; adapter_layer; bach2022promptsource). Information flow is conditionally routed to a subset of these modules, where it is then aggregated. This design facilitates the positive transfer and systematic generalization (pfeiffer2023modular). Recently, Poly and MHR were designed to tackle diverse tasks by leveraging different combinations of latent experts (poly; mhr). Given $|\mathcal{T}|$ tasks, there are only $|\mathcal{K}|<|\mathcal{T}|$ experts are trained, we treat these $|\mathcal{K}|$ experts as latent experts, and each task-specific adapter can be obtained by a linear combination of latent experts. These approaches allow for the recombination of previously learned experts to solve test tasks more effectively, showcasing a significant advancement in task adaptability and learning efficiency. During both the multi-task pre-trained and fine-tuning, Poly/MHR concurrently optimizes the parameters of the expert inventory and a task-experts routing function, where each task-specific expert expressed by a linear combination of the activated experts. To this end, a natural question is: can we craft a non-linear function to merge experts?
Figure 1: LoRA reparamarized by an entangled tensor with a rank = 2 and order = 3. Different granularity of modules induces two routing approaches. TensorPoly-I will select the rank of the tensor while TensorPoly-II select the finer-grained order of the entangled tensor.
To answer this question, we propose a variant Poly model: TensorPoly, which mixture experts using tensor product (smolensky1990tensor). The tensor product $V\times W$ of two vector spaces $\mathcal{V}$ and $\mathcal{W}$ is a vector space associated with a bilinear map $V\times W\rightarrow V\times W$3.2). This process enables the capturing of higher-order interactions and structural relationships between the input spaces (panahi2019word2ket; kye2023compositions; gan2022morphte). As depicted in Figure 1, we have reparameterized LoRA (lora) adapters by employing an "entangled" tensor structure. Consequently, the traditional training matrix $M\in\mathbb{R}^{d\times r}$ in LoRA is reparameterized into a more finer-grained tensor $\mathcal{L}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil\times R)}$, named as TLoRA (4.1). This reparameterization allows for a more nuanced manipulation of the model’s parameters, facilitating a more precise and efficient adapter process. The entangled tensor configuration introduces two critical hyper-parameters: the tensor rank (R) and the tensor order (N). Leveraging these parameters, we have developed two distinct routing functions designed to select modules with varying levels of granularity. As depicted in Figure 1, TensorPoly-I employs a routing mechanism that assigns distribution scores to different tensor ranks, facilitating the selection of modules based on their rank granularity. Further advancing this concept, we propose a more refined routing function, TensorPoly-II, which targets even finer-grained tensors as activated modules. Each module is associated with a specific order of the entangled tensor. Once modules are selected via the routing function, they are aggregated through a tensor product operation, enabling a sophisticated and dynamic assembly of modular skills.
We evaluate our methods against a series of competitive baselines for few-shot task adaptations benchmark (t0_bench). Our experiments reveal several key insights:
  1. 1.
    Modular language models such as Poly and TensorPoly frameworks, consistently outperform traditional PEFT approaches LoRA and TLoRA, underscoring the effectiveness of our routing function in facilitating positive transfer across multi-task environments.
  2. 2.
    TensorPoly-I demonstrates competitive results as Poly, while simultaneously reducing the number of training and adaptation parameters required. This efficiency gain highlights the benefits of our tensorized module approach in achieving high performance with lower parameter overhead.
  3. 3.
    A comparative analysis between TensorPoly-I and TensorPoly-II indicates that the latter’s finer-grained routing mechanism does not contribute to improved performance in tensor product routing scenarios. This outcome suggests that while granularity in module selection is valuable (mhr), there is a complexity threshold beyond which additional granularity may not yield further benefits.
In summary, our contributions are as follows:
  • We propose a novel PEFT approach TLoRA that achieves competitive results while only use $\mathcal{O}(N\times r\times\lceil\sqrt[N]{d}\rceil\times R)$ training parameters compared to LoRA $\mathcal{O}(d\times r)$, highlighting our approach’s high parameter efficiency.
  • We propose a novel modular language model TensorPoly using tensorized modules TLoRA, designed to enhance positive transfer and generalization across multi-task scenarios.
  • The findings from our T0 benchmark demonstrate that our modular language model significantly enhances the performance of a single module, underscoring the critical role of routing in scenarios involving multiple tasks.
2 Related Work
PEFT Parameter efficient fine-tuning (PEFT) methods facilitate efficient adaptation of LLMs without updating all the training parameters, thereby reducing the memory and computation cost (zhang2023llama). One kind of PEFT approach focuses on adding modules to LLMs, and only these small modules will be trained while the backbone model is kept frozen and shared across tasks. For example, Adapter Tuning inserts small neural modules (adapters) between the layers of the basic model (adapter_layer), whereas Prefix Tuning and Prompt Tuning add tunable vectors to the input or hidden layer of the base model (li2021prefix; lester2021power). Another kind of research model is the incremental update of the pre-trained weights in a parameter-efficient way, without modifying the model architecture. Bit kit fixes all training parameters and only finetunes the additive bias term (zaken2021bitfit). Diff Pruning learns a task-specific “diff vector” that extends the original pre-trained parameters. As the number of tasks increases, Diff Pruning only requires storing only a small diff vector for each task (guo2020parameter). The seminal paper lora proposes a method named LoRA that parameterizes incremental weights $\Delta$ as a low-rank matrix by the product of the down projector matrix and up projector matrix. LoRA achieves comparable or even better performance than full fine-tuning (lora). zhang2023adaptive demonstrates that weight matrices in the top layers are more important than those in the bottom layers. They propose a new method: Adaptive Low Rank Adaptation (AdaLoRA) that dynamically allocates the parameter budget among weight matrices during LoRA-like finetuning (zhang2023adaptive). AdaLoRA adjusts the rank of incremental matrices for different layers. liu2024dora introduces a new approach DoRA to investigate the inherent differences between full finetuning and LoRA. In our approach, TLoRA reparameterize LoRA using tensor products, utilizing the finer-grained tensors as modules.
MTL To share the information across multiple tasks, AdapterSoup chronopoulou2023adaptersoup trains each adapter for each domain, then it performs weight-space averaging of adapters trained on different domains. (huang2023lorahub) introduce LoRAhub to aggregate the LoRA modules trained on diverse tasks. They first train a group of LoRA modules which are specialized in each task, then randomly select a subset of modules, and finally learn a set of weights to combine these LoRA models using gradient-free optimization. AdapterFusion (pfeiffer2020adapterfusion) proposes a two-stage algorithm that leverages knowledge from multiple tasks. Similarly as in LoRAhub, a group of task-specific adapters learn to encapsulate the task-specific information, and in a second stage, a fusion layer combine the trained adapters. poly introduces a variable-size module routing mechanism, Poly, predicated on the notion that each task correlates with a specific subset of latent skills drawn from a comprehensive inventory of modules. Concurrently, both the modules and the routing function are trained, facilitating a more dynamic and task-specific approach to model learning. Building upon Poly, (mhr) introduces a finer-grained multi-head routing function MHR where the experimental findings underscore the significance of the routing function during the pre-training phase. Remarkably, by solely fine-tuning the routing function, MHR attains competitive results, demonstrating exceptional parameter efficiency.
Model Merging Model merging can be seen as a parameter level aggregation, combining multiple models into one single model in parameter space without access to data (matena2022merging). The simplest operation of merging is averaging the weights of different models. wortsman2022model shows that averaging the weights of multiple models finetuned with different hyperparameter configurations often improves accuracy and robustness. matena2022merging takes the aggregation function maximizing the joint likelihood of the posteriors of the model’s parameters and proposes a “Fisher merging” technique. yadav2023ties proposes a TIES-Merging method to address interference due to redundant parameter values and disagreement on the sign of a given parameter’s values across models. ilharco2022editing edits models with task arithmetic operations. jin2022dataless propose a dataless knowledge fusion method (Regression Mean) that merges models in their parameter space.
3 Background
In multiple transfer learning tasks , we define a set of tasks as $\mathcal{T}=\{\mathcal{T}_{1},...,\mathcal{T}_{|\mathcal{T}|}\}$. This set is divided into two subsets train $\mathcal{T}_{train}$ and test $\mathcal{T}_{test}$. The goal of multi-task transfer learning is to apply the knowledge from the training tasks $\mathcal{T}_{train}$ to the test tasks within $\mathcal{T}_{test}$. This process involves two main phases. Building on top of a foundation model, the first phase consists of multi-task pre-training using the dataset in tasks $\mathcal{T}_{train}$. The second consists of a few-shot adaptation, where the learned adapters are fine-tuned independently on each test task in $\mathcal{T}_{test}$. We follow the procedure from (t5) and formulate each task as a text-to-text problem, enabling standard maximum-likelihood training with teacher forcing (bengio2015scheduled) and a cross-entropy loss.
3.1 Module: LoRA
LoRA are recently proposed adapter architecture that achieve a competitive balance between performance and parameter efficiency (lora; mahabadi2021parameter). For each linear transformation corresponding to the query ($q$), key ($k$), value ($v$) and output ($o$) of the self-attention layers, LoRA modifies the base model parameters as follows:
$h=W_{0}x+s\cdot A(B)^{\top}x$ (LoRA)
where $W_{0}$ are the (frozen) weights of the pre-trained model (e.g. T5 (t5)). $A,B\in\mathbb{R}^{d\times r}$ are low-rank learnable parameters and $s\geq 1$ is a tunable scalar hyperparameter. We schematize LoRA in Figure 2.
3.2 Tensor, Tensor Product, Entangled Tensor
Tensor.
The tensor $\mathcal{A}$ is a multidimensional array of elements (called components) of $\mathbb{R}$, each denoted by its integer coordinates in the array; e.g., for a two-dimensional array, the component at position $i,j\in\mathbb{N}$ is denoted $A_{i,j}$. The order of a tensor is how many indices it has (e.g., a vector $v$ is a first-order tensor, a matrix $M$ is a second-order tensor, etc.)
Tensor Product.
The tensor product $V\otimes W$ of two vector spaces $\mathcal{V}$ and $\mathcal{W}$ is a vector space to which is associated a bilinear map $V\times W\rightarrow V\otimes W$ that maps a pair (v, w), $v\in V,w\in W$ to an element of $V\otimes W$ denoted $v\otimes w$. We can create tensor product spaces by more than one application of tensor product, $\mathcal{H}=\mathcal{U}\otimes\mathcal{V}\otimes\mathcal{W}$, with arbitrary bracketing, since tensor product is associative. Tensor product space of the form is said to have tensor order of $n$.
$\bigotimes_{j=1}^{n}\mathcal{H}_{j}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes...\otimes\mathcal{H}_{n}$ (1)
Entangled Tensor.
The $n$-order tensor product space $\otimes^{n}_{j=1}\mathcal{H}_{j}$ consists of vectors of the form $v=\otimes^{n}_{j=1}v_{j}$, where $v_{j}\in\mathcal{H}_{j}$, are called simple tensor. Vectors need to be represented as the sum of multiple simple tensors called entangled tensors:
$\sum_{k=1}^{r}\bigotimes_{j=1}^{n}\mathcal{H}_{j}=\sum^{r}_{k=1}\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes...\otimes\mathcal{H}_{n}$ (2)
where tensor rank $r$ is is the smallest number of simple tensors that sum up to $v$. For example, $\frac{\psi_{0}\otimes\phi_{0}+\psi_{1}\otimes\phi_{1}}{\sqrt{2}}$ is a tensor of rank 2.
3.3 Tensorized Vector using Engangled Tensor
Any training parameters $v\in\mathbb{R}^{d}$ can be expressed as an entangled tensor of rank $r$ and order $n$ by: $v=\sum\limits_{k=1}^{r}\bigotimes_{j=1}^{n}v_{jk}$. Here, $v_{jk}\in\mathbb{R}^{q}$, yielding a resultant vector $v$ of dimension $p=q^{n}$. Its storage requirements are efficiently managed, consuming only $rnq=O(rq\log p/q)$. If $q^{n}>d$, the excess part of the generated vector will be cut off. The number of vector parameters can be reduced from $d$ to $rn\sqrt{d}$. For example, when $d=512$, $q=8$, $n=3$ and $r=2$, the number of parameters of a vector can be reduced from 512 to 48.
4 Methods: TensorPoly
Figure 2: TensorPoly-I and TensorPoly-II. We illustrate how to reparametrized the LoRA matrix $\mathbb{R}^{625\times 5}$ with 4 tensor $\mathcal{A}\in\mathbb{R}^{3\times 5\times 5}$. In this case, the tensor rank $R=3$, tensor order $N=4$. For TensorPoly-I, the routing function $Z$ is designed to select which rank of the entangled tensor is activated for a given task. Conversely, TensorPoly-II introduces a more granular control by selecting tensor rank and tensor order. The Blue color indicates a non-trainable component and reddish color indicates a trainable component
In this section, we introduced the Poly model. This section delves into two variant models, TensorPoly-I and TensorPoly-II. TensorPoly-I employs a routing mechanism that assigns distribution scores to tensor ranks while TensorPoly-II assigns distribution scores to the finer-grained tensor order.
4.1 TLoRA
TLoRA reparamizes the LoRA using the tensor product.
$\begin{split}h&=W_{0}x+s\cdot A(B)^{\top}x\\ &=W_{0}x+s\cdot\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{A}_{i,k}\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\mathcal{B}_{i,k}x\end{split}$ (TLoRA)
where $R$ is the rank, $N$ is the order of entangled tensor respectively.
4.2 Polytropon (Poly)
Poly/ MHR addresses the multi-task problem by softly sharing latent experts across tasks. Each Poly layer contains 1) an inventory of latent experts =$\mathcal{M}=\{\phi_{1},\ldots,\phi_{m}\}$ with $|\mathcal{M}|\ll|\mathcal{T}|$; 2) a routing function $r(\cdot)$ that chooses which subset of the experts to combine for each task.
Each module corresponds to a LoRA adapter, where $\phi_{i}$ are its associated parameters $A^{(i)},B^{(i)}\in\mathbb{R}^{d\times r}$. $r(\cdot)$ is implemented as a task–module routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{M}|}$. $z_{\tau}=Z_{\tau,:}\in\mathbb{R}^{|\mathcal{M}|}$ is a routing vector of task ${\mathcal{T}_{\tau}}$, with cell $Z_{\tau,j}$ being the probability logits of using module $\phi_{j}$ for task $\mathcal{T}_{\tau}$ in the current layer. Differently from mixture-of-experts (fedus2021switch), which perform token-level top-$k$ routing, $Z$ converges to a binary matrix, defining a soft partition over modules. This is achieved by using a Gumbel-sigmoid distribution (jang2016categorical) during training, with $\hat{Z}_{\tau,j}\sim\texttt{Gumbel}(Z_{\tau,j})$. At each forward pass, Poly can be defined as :
$A^{\tau}=\sum_{i}\alpha_{i}A^{(i)};\,B^{\tau}=\sum_{i}\alpha_{i}B^{(i)}$ (Poly)
where $\alpha_{i}=\frac{\hat{Z}_{\tau,i}}{\sum_{j}\hat{Z}_{\tau,j}}$ , and $A^{(i)},B^{(i)},A^{\tau},B^{\tau}\in\mathbb{R}^{d\times r}$. We normalize the mixing coefficients $\hat{Z}_{\tau,i}$ for each task to ensure that the number of active modules does not affect the norm of $A^{\tau},B^{\tau}$. Overall, this approach enables different subsets of modules to be activated for the current layer and combined in a task-specific way. Following TLoRA, the output of the Poly layer is added to the output of the original layer of the frozen backbone: $h=W_{0}x+sA^{\tau}(B^{\tau})^{\top}x.$
4.3 TensorPoly-I
As illustrated in Figure 2, the TensorPoly model innovatively incorporates the Tensorized Low-Rank Adaptation (TLoRA) as its core module. This approach adopts finer-grained tensors for training parameters, thereby enabling a more nuanced parameter manipulation. The routing matrix $Z\in\mathbb{R}^{|\mathcal{T}|\times R}$ plays a pivotal role in this framework, determining which rank within the entangled tensor is to be activated for a given task. Upon activation of the selected rank, the model performs a linear combination of the selected tensor, weighted by a factor $\alpha$:
$A^{\tau}=\sum\limits_{k=1}^{R}\alpha\bigotimes_{i=1}^{N}\mathcal{A}_{i,k};B^{\tau}=\sum\limits_{k=1}^{R}\alpha\bigotimes_{i=1}^{N}\mathcal{B}_{i,k}$ (TensorPoly-I)
where $\alpha=\frac{\hat{z}_{\tau,i}}{\sum_{j}\hat{z}_{\tau,j}}$. $\mathcal{A}_{i,k}\in\mathbb{R}^{N\times r\times\lceil\sqrt[N]{d}\rceil}$ is a third-order tensor.
4.4 TensorPoly-II
In TensorPoly-II, we implement a fine-grained routing function that selects both the tensor rank and tensor order, enabling precise control over the model’s adaptation to specific tasks. This sophisticated routing mechanism facilitates the selection of finer-grained tensor elements, which are then aggregated through a tensor product operation followed by a linear combination. The routing itself is conceptualized as a third-order tensor $Z\in\mathbb{R}^{|\mathcal{T}|\times R\times N}$, which offers an unprecedented level of granularity in directing the model’s focus across different ranks and orders of the tensor space.
$A^{\tau}=\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\alpha\mathcal{A}_{i,k};B^{\tau}=\sum\limits_{k=1}^{R}\bigotimes_{i=1}^{N}\alpha\mathcal{B}_{i,k}$ (TensorPoly-II)
Each unit of $Z$ is $\alpha=\frac{\hat{z}_{\tau,i,k}}{\sum_{j}\hat{z}_{\tau,j,k}}$, which will be routed to the specific order and rank in the entangled tensor.
5 Experiments
Model Natural Language Inference Sentence Completion Co-reference WSD ACC Param
RTE CB ANLI1 ANLI2 ANLI3 COPA H-SWAG Story WSC Wino WiC
Baselines (w/o pretrain)
FullFT 79.8 87.5 46.6 41.3 40.0 81.0 46.4 93.8 65.4 56.5 57.7 63.3 3B
BitFit (with LN) 72.2 57.1 36.5 35.3 36.6 75.0 29.5 88.6 61.5 56.6 51.7 54.6 1.3M
LayerNorm 71.8 57.1 36.5 35.1 36.3 76.0 29.6 88.7 63.5 49.4 52.2 54.2 250K
Adapter 76.2 87.5 45.1 40.4 35.3 84.0 41.9 91.7 65.4 54.7 55.5 61.6 12.9M
Compacter 75.8 82.1 40.8 37.4 35.8 84.0 46.4 93.5 64.4 55.5 55.2 61.0 807K
Compacter++ 76.9 82.1 41.7 38.3 36.9 86.0 46.3 93.5 65.4 55.1 54.1 61.5 540K
Prompt(10) 52.7 66.1 34.2 33.5 33.5 67.0 29.9 84.2 54.8 51.9 51.6 50.9 41K
Prompt(100) 48.0 53.6 33.4 33.8 33.3 60.0 26.8 74.0 60.6 51.1 50.0 47.7 409K
Prefix tuning 68.6 84.0 43.3 37.5 36.5 71.0 42.1 90.2 56.7 52.0 54.2 57.8 576K
FishMask (0.2%) 76.9 83.9 43.7 39.7 37.2 82.0 44.1 94.2 63.5 54.5 52.5 61.1 6M
FishMask (0.02%) 75.5 76.8 39.9 38.1 36.2 84.0 38.2 93.6 61.5 53.9 53.5 59.2 600K
SAID 69.0 80.4 40.4 35.4 35.5 77.0 36.7 89.3 61.5 52.7 55.0 57.5 500K
SAID 66.1 83.9 41.3 38.5 35.8 76.0 38.3 89.7 55.8 50.9 55.3 57.4 20K
LoRA 78.3 85.7 45.1 41.0 39.5 88.0 47.1 93.6 60.6 56.8 55.2 62.8 9.1M
$\mathrm{(IA)}^{3}$ 78.0 87.5 48.6 40.8 40.8 87.0 49.4 94.7 68.3 59.8 56.0 64.6 540K
DoRA
w/ pretrain
LoRA 81.9 89.3 41.2 40.3 41.3 93.7 59.8 96.2 66.0 67.9 56.8 66.8 2.2M
$\mathrm{(IA)}^{3}$ 82.2 89.9 45.8 41.6 41.2 91.7 53.4 94.2 70.8 63.3 53.9 66.2 540K
DoRA
TLoRA 80.7 90.5 39.9 40.9 41.2 93.0 54.4 95.3 66.3 67.4 57.3 66.1 1.4M
Poly 84.7 89.3 46.0 42.8 42.7 93.0 63.3 96.6 68.9 70.1 59.9 68.8 17M
MHR 85.2 90.5 44.7 42.3 42.8 94.7 63.3 96.7 70.5 70.6 59.8 69.2 17M
TensorPoly-I 85.2 91.7 45.0 42.5 42.5 96.7 63.1 96.6 68.6 69.8 60.6 69.3 27.8M
TensorPoly-II 84.7 90.5 44.4 41.0 42.0 94.3 58.7 95.7 67.6 68.7 59.9 68.0 33.2M
Table 1: Results on the T0 few-shot benchmark. All the results in our implementation are the median score of 3 seeds [0,1024,42]. For all the baseline scores, we report the results from tfew. The bold is the best score. The underline is the second best in the same training setting.
Our experiments aim to answer the following questions: 1) How are different PEFT approaches performed? 2) Is routing important in multi-task scenarios? 3) Can non-linear merging methods in TensorPoly generalize well?
5.1 Backbone, Datasets and Evaluation
Backbone
To ensure our model retains high performance with a limited number of labeled examples after fine-tuning, it is crucial to select an appropriate pre-trained model as the backbone. To facilitate a fair comparison with baseline methodologies, we have chosen the T0 model, consistent with the approach described in the IA3 paper by tfew. T0 was created by finetuning the T5 model on a multi-task mixture of datasets (t5). Each dataset is associated with multiple prompt templates that are used to format the example to (input, target) pairs. Examples in the datasets to train the T0 were prompted by applying the prompt templates from the Public Pool of Prompts (P3) (bach2022promptsource). As a result, all the different types of natural language tasks can trained with sequence-to-sequence structure (t0_bench).
Datasets
To evaluate the generalization capabilities of our models, we adopt the same benchmarking strategy as (tfew), utilizing a subset of tasks designated as held-out from the multitask training. This benchmark encompasses a diverse array of tasks, including sentence completion (COPA (roemmele2011choice), H-SWAG(zellers2019hellaswag) and Story Cloze (sharma2018tackling) datasets), natural language inference (ANLI (nie2019adversarial), CB (de2019commitmentbank) and RTE (dagan2005pascal)), coreference resolution (WSC (levesque2012winograd), Winogrande(sakaguchi2021winogrande)), and word sense disambiguation (WIC (pilehvar2018wic)). For each task, our evaluation strategy involves constructing sets of five few-shot training examples, which are generated by sampling subsets from each dataset using different seeds. We then report the median performance. It is note that the prompt examples from each dataset using the prompt templates from P3 (bach2022promptsource), using a randomly-sampled prompt template for each example.
Evaluation
For the evaluation of our models, we employ the rank classification methodology as outlined by the tfew study. This approach involves ranking the model’s log-probabilities for all possible label strings associated with each task. The model’s prediction is deemed correct if the label string with the highest log-probability ranking corresponds to the correct answer. This method allows for a nuanced assessment of the model’s predictive accuracy by examining its ability to prioritize the correct label over others based on their calculated log-probabilities, offering a precise measure of its understanding and processing of the task at hand.
5.2 Baselines
In our comparative analysis, we initially set the benchmark by evaluating the performance of the TLoRA model against the traditional full fine-tuning approach, referred to as FullFT. In the FullFT scenario, we do not freeze any parameters of the pre-trained model, nor do we insert any adapters, allowing for a comprehensive update of the model’s parameters during fine-tuning. Subsequently, we contrast our method against a suite of established parameter-efficient fine-tuning (PEFT) baselines to study the efficiency and effectiveness of each in terms of training parameter utilization. These baselines include:
  • Adapter, as introduced by houlsby2019parameter, which involves inserting trainable layers while keeping the pre-trained model’s parameters fixed.
  • BitFit by zaken2021bitfit, which only fine-tunes the bias terms within the model.
  • LoRA proposed by lora, adjusting the low-rank adaptations of the weight matrices.
  • Compacter and Compacter++ by karimi2021compacter, which extend the adapter methodology with compact and efficient training strategies.
  • Prompt tuning (lester2021power) and Prefix tuning li2021prefix add some tunable vectors to the input or hidden layer of the base model.
  • FishMask by sung2021training, identifying and training a subset of parameters.
  • Intrinsic SAID as described by aghajanyan2020intrinsic, focusing on intrinsic sparse activations.
  • IA3 (tfew), emphasizing adaptability and efficiency.
We undertake a comparative analysis between routing approaches TensorPoly, Poly and employing a single expert mechanism without routing, specifically TLoRA and LoRA. This comparison aims to evaluate the impact of routing techniques on model performance and efficiency. By contrasting these models, we seek to understand how the dynamic allocation of tasks to specific experts in TensorPoly and Poly compares to the single expert models without shared information across tasks.
5.3 Results
Method Pre-Training Fine-Tuning
FullFT $d\times d$ $d\times d$
LoRA $2\times d\times r$ $2\times d\times r$
TLoRA $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$
Poly $2\times d\times r\times|S|+|\mathcal{T}|\times|S|$ $2\times d\times r\times|S|+|S|$
TensorPoly-I $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R+|\mathcal{T}|\times R$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ + $R$
TensorPoly-II $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R+\mathcal{T}\times R\times N$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ + $R\times N$
Table 2: Number of parameters (per layer) used for each method. $d$ is the input and output dimension of the training parameters. We assume they are identical. $r$ is the rank in the LoRA, where $r\ll d$. $N$ and $R$ are the order and rank of entangled tensor respectively. $S$ is the number of modules in Poly.
Table. 1 presents the mean downstream accuracy for 11 held-out tasks in the T0 benchmark. Notably, we did not train all these PEFT approaches, all the results are reported from (tfew). When evaluating the performance of various PEFT approaches against single expert performances, it is observed that many PEFT strategies achieve similar outcomes while utilizing a significantly smaller subset of training parameters compared to the FULLFT (Full Fine-Tuning) method. Additionally, the implementation of Low-Rank Adaptation (LoRA) within our training framework further illustrates the efficiency of these methods. Specifically, TLoRA achieves a competitive score of 66.1, closely trailing the original LoRA’s score of 66.8, while requiring only about 60% of the training parameters used by LoRA. This demonstrates that TLoRA not only matches the effectiveness of LoRA in terms of performance but also surpasses it in terms of parameter efficiency.
In our analysis, we initially contrast the modular model against the dense model, followed by a comparison of routing-based approaches with a single adapter strategy. Within this context, we utilize both LoRA and TLoRA as baselines for these routing techniques. According to the results presented in Table 1, the Poly model demonstrates superior performance over LoRA by a margin of 2.0 points. Moreover, TensorPoly exhibits an improvement of 3.2 points over the base TLoRA model, underscoring the efficacy of routing in enhancing multi-task generalization within the realm of multi-task transfer learning.
When evaluating various Poly variants, TensorPoly-I stands out by not only surpassing recent state-of-the-art achievements but also by outperforming TensorPoly-II, despite the latter employing a more granular routing function. This finding is particularly noteworthy, as it suggests that the increased specificity of the routing function in TensorPoly-II does not necessarily translate to superior performance. We will discuss this in Section 5.5.
5.3.1 Rank and Order Analysis
Figure 3: Explore the tensor Rank in the Entangled Tensor. We set order $N=2$ in this setting.
The rank of an entangled tensor serves as a measure of its capacity, denoting the fewest number of rank-one tensors required for its original tensor (§3.2). We focus on examining the relationship between the rank of tensor and the associated training parameters. Table 2 illustrates how the tensor rank $R$ and tensor order $N$ impact the selection of training and fine-tuning parameters. Illustrated in Figure 3, when the tensor rank is set to 1, it corresponds to the original TLoRA configuration. An increase in tensor rank necessitates a larger number of training parameters. Concomitantly, there is a notable enhancement in performance. This progression underscores the direct correlation between the tensor rank and the model’s capability.
Then tensor order $N$, correlates with the granularity of training parameters; a tensor of order 4 yields a finer-grained module compared to one of order 2. Our research examines the impact of varying tensor orders on performance outcomes. For simplicity, we constrain our analysis to tensors of order 2 and 4. Our results demonstrate that a tensor of order 2 outperforms one of order 4 by a margin of 1.7 for the TensorPoly-I. For TensorPoly-II, an order-2 tensor exceeds the performance of an order-4 tensor by 4.8, suggesting a balance must be struck between the size of experts and its efficacy. While higher-order tensors may conserve training parameters, this comes at the cost of diminished performance.
Model Adaptation Params ACC
LoRA[r=1] 2.2M 66.8
TLoRA [r=16] 1.4M 66.1
Poly[r=1] 17 M 68.8
TensorPoly-I[N=2] 12.2M 68.7
TensorPoly-I[N=4] 4.3M 66.9
TensorPoly-II[N=2] 13.3M 67.9
TensorPoly-II[N=4] 7.6M 63.1
Table 3: Results on T0 dataset, with T0-3B. The Adaptation parameter is the number of parameters in the finetuning process.
Model Adaptation Params ACC
LoRA[r=1] 2.2M 66.8
TLoRA [r=16] 1.4M 66.1
Poly-$\mu$[r=1] 2.2M 68.8
MHR-$\mu$ 2.2M 68.6
TensorPoly-I-$\mu$[N=2] 1.4M 69.0
TensorPoly-II-$\mu$[N=2] 1.4M 68.2
Table 4: Finetuning the modules only. The Adaptation parameter is the number of parameters in the finetuning process.
5.4 Routing Analysis
(a) Multi-choice QA
(b) Summarization
(c) Closed-Book QA
Figure 4: Distribution of experts weights in the T0 model. X-axis is the 32 experts in each layer (For Q,K,V,O, each corresponds to 8 experts). Y-axis denotes the 24 transformer layers in T0 model. We show routing distributions for different kinds of tasks.
In the paper referenced as mhr, the MHR study demonstrates that fine-tuning solely the routing function can yield competitive outcomes. This insight provides a valuable perspective for our investigation into various routing strategies within the TensorPoly framework. In line with this approach, we focus on fine-tuning exclusively the routing function during the few-shot adaptation process, indicated by the notation -$z$. This methodological choice allows us to isolate the impact of the routing function’s optimization on the overall performance of the TensorPoly model, thereby offering a clearer understanding of how dynamic routing contributes to the adaptability and efficiency of the model in few-shot learning scenarios.
As detailed in Table 5, an initial observation reveals that the routing parameters necessitate a remarkably small number of training parameters. The performance metrics of Poly-$z$, TensorPoly-I-$z$, and TensorPoly-II-$z$ lag behind their counterparts where both the modules and the routing function undergo fine-tuning. This disparity highlights the critical role that modules play in the fine-tuning process for TensorPoly. The findings suggest that while optimizing the routing function alone can contribute extreme parameter efficiency, the integration of module fine-tuning is indispensable for achieving the best possible performance. This underlines the synergy between modules and routing functions in the TensorPoly architecture.
Model Adaptation Params ACC $\Delta$
Poly-$z$ 3.5k 65.4 -3.4
MHR-$z$ 220K 68.3 -0.9
TensorPoly-I-$z$[N=2] 3.5k 64.3 -4.4
TensorPoly-I-$z$[N=4] 3.5k 62.7 -0.4
TensorPoly-II-$z$[N=2] 6.9k 62.5 -5.4
TensorPoly-II-$z$[N=4] 13.8k 60.3 -2.8
Table 5: Results on T0 dataset, with T0-3B. We only fine-tuning the routing function.
5.5 Discussion
In modular language models, tensor product routing is poised to construct more intricate patterns. This capability arises from its method of aggregating activated modules through tensor products, as opposed to the conventional linear routing approach. However, our experimental results diverged from theoretical expectations, revealing that tensor product routing did not contribute to improved final performance in our models. This discrepancy prompts a future line of inquiry: we plan to explore whether there exist specific benchmarks or conditions under which tensor product routing could demonstrate its purported benefits. Identifying such scenarios will be crucial for harnessing the potential advantages of tensor product routing in modular language models.
In the current methodology, each tensor-based module is not specialized towards any particular domain. In the future, we intend to explore a more tailored training strategy. This will involve dedicating each tensor to a specific domain and subsequently aggregating these domain-specialized tensors using tensor product operations. Our objective is to assess whether this domain-specific aggregation approach can yield superior generalization capabilities.
6 Conclusion
We introduce a novel modular language model named TensorPoly. This model incorporates tensorized modules, specifically TLoRA, to significantly reduce the number of training parameters required by the traditional LoRA approach. We employ two distinct strategies for aggregating activated modules: a linear routing function, referred to as TensorPoly-I, and a more finely trained tensor product routing, named TensorPoly-II. Our evaluation across various multi-task learning scenarios reveals that modular language models, such as TensorPoly, surpass the performance of single-adapter models. This underscores the importance of sharing task information through a routing function in multi-task learning contexts. Notably, TensorPoly-I achieves state-of-the-art results, highlighting the effectiveness of the TensorPoly framework. However, TensorPoly-II does not outperform TensorPoly-I in our experimental settings, suggesting areas for further investigation in future research.
Broader Impact Statement
The research reported in this paper proposes a novel type of language model, and is primarily a theoretical contribution accompanied by experiments to show the practical usefulness on the model, especially in mitigation of negative inference in multi-task learning. Thus, the impact is primarily in generic modelling and empirical performance in situations where existing approaches (e.g., LoRAs) are used, but does not per se create new use scenarios, or raise new concerns, ethical or otherwise, beyond those already present for LoRAs.
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors. Only add this information once your submission is accepted and deanonymized.
Acknowledgments
Anonymized for double-blind review.
Appendix A Appendix
A.1 FLOP Analysis
TLoRA aims to reduce the number of training parameters by parameterizing the original LoRA architecture, but it does not alter the computational process of a LoRA. Thus, TLoRA might conceivably incur additional computational overhead compared to a standard LoRA. To understand this, we analyze the floating-point operations (FLOPs) involved in TLoRA.
Consider the tensor product of a vector $a$ of size $n\times 1$ with a vector $b$ of size $1\times n$, resulting in a matrix $C$ of size $n\times n$. The product operation involves $n^{2}$ multiplications due to the $n$ elements in each of the vectors $a$ and $b$. As described in Table 2 and shown in Figure 2, the additional computational effort in TLoRA stems from the tensor product operations, dictated by the order of the tensor. Specifically, for parameterizing with a dimension $d$, rank $r$, and tensor rank $R$, the number of extra computations required is approximately $d\times r\times R$. There may add some time consumption analysis
N=1,R=1
N=1,R=2
N=1,R=3
N=1,R=4
N=1,R=5
N=1,R=6
N=1,R=7
N=1,R=8
N=2,R=1
N=1,R=2
N=1,R=3
N=1,R=4
N=1,R=5
N=1,R=6
N=1,R=7
N=1,R=8
Figure 5: Weights distributions across different experts. We set N=2, R=8.